winning symbols under the cap of 30% of all cola bottles, if you buy 2 pack of cola what is the probability that you win something
Answer #1

P(winning when you buy a cola) = 0.3

P(not winning when you buy a cola) = 1 - 0.3 = 0.7

P(winning something when you buy 2 cola) = 1 - P(not winning from both colas)

= 1 - 0.7x0.7

= 0.51

P.S: Please comment if a pack of cola means more number of soda bottles, like 12 bottles in a pack. The answer will vary in that case. For example, if a pack is having 12 colas, P(winning something) = 1 - 0.712x2

= 0.9998